8x-15=5x
5x 5x
13x=-15
x=13/15

Answer by Alan3354(69443)

8x-15=5x
5x 5x
13x=-15 *****
x=13/15
=================
8x-15=5x
5x 5x
--------------- Subtract
3x-15 = 0
Add 15
3x = 15
x = 5